OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 45, what is the equation of the line? 4y − 5x = 5 y − 4x = 5 5y + 4x = 5 5y − 4x = 5 @Insert_Name_Here

OpenStudy (anonymous):

y-axis(0,1)= 4x +5y=5

OpenStudy (anonymous):

c? thats wrong

OpenStudy (anonymous):

oh sorry! that plus was suppose to be a minus 0.o

OpenStudy (anonymous):

I got it wrong ): Im gonna retake it

OpenStudy (anonymous):

Im sorry for the mistake.. :'(

OpenStudy (anonymous):

If a line with a slope of -2 crosses the y-axis at (0, 3), what is the equation of the line?

OpenStudy (anonymous):

y − 2x = 3 -2y + 3x = 1 y + 3x = 2 y + 2x = 3

OpenStudy (anonymous):

the ????? is a - (subraction/minus)

OpenStudy (anonymous):

@Insert_Name_Here

OpenStudy (anonymous):

its gonna be a negative #?

OpenStudy (anonymous):

Negative?

OpenStudy (anonymous):

No IS it gonna be a negative?

OpenStudy (anonymous):

The answer? idk

OpenStudy (anonymous):

I think its either A or D

OpenStudy (anonymous):

It was D

OpenStudy (anonymous):

i thought so

OpenStudy (anonymous):

I didnt want to give you the wrong answer

OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 45, what is the equation of the line?

OpenStudy (anonymous):

4y − 5x = 5 y − 4x = 5 5y + 4x = 5 5y − 4x = 5

OpenStudy (anonymous):

@Insert_Name_Here

OpenStudy (anonymous):

Hmmmmm.. so its not C or D

OpenStudy (anonymous):

So its either A or B right?

OpenStudy (anonymous):

Im leaning towards A

OpenStudy (anonymous):

wbu

OpenStudy (anonymous):

@Radelp

OpenStudy (anonymous):

the slope is 45? or 4/5?

OpenStudy (anonymous):

Wow ok

OpenStudy (anonymous):

I think it is A aswell

OpenStudy (anonymous):

It wasnt showing your replys @Insert_Name_Here

OpenStudy (anonymous):

I SAID IT WAS A!!

OpenStudy (anonymous):

ugh nevermind

OpenStudy (anonymous):

I thought it was C

OpenStudy (anonymous):

I didnt see it until now SORRY

OpenStudy (anonymous):

lol be patient man and c is actualy right, sorry i saw something wrong ur good @savannahPaige18

OpenStudy (anonymous):

C was wrong tho

OpenStudy (anonymous):

the values are all positve^

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

CRAP sorry... :'(

OpenStudy (anonymous):

maybe it is A i may retake it. Still got more questions

OpenStudy (anonymous):

Sorry @Insert_Name_Here ans @SavannahPaige18

OpenStudy (anonymous):

WTF?! Buh she tried C and it was wrong

OpenStudy (anonymous):

U know what.....screw it

OpenStudy (anonymous):

\(\bbox[9pt,#f9bec7,border:7pt solid #33cccc]{\color{white}{\Huge\cal\dagger I'm~Out \dagger}}\)

OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 4x + 5y = -9.

OpenStudy (anonymous):

slope = 45 and y-intercept = −95 slope = −45 and y-intercept = 95 slope = −45 and y-intercept = −95 slope = 45 and y-intercept = 95

OpenStudy (anonymous):

@Insert_Name_Here @Radelp

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

i tagged you in a new question